TANGLE(1) TANGLE(1) NAME tangle−translateWEBtoPascal SYNOPSIS tanglewebfile[.web][changefile[.ch]] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. ThetangleprogramconvertsaWEBsourcedocumentintoaPascalprogramthatmaybecompiledintheusualwaywiththeon-linePascalcompiler(e.g.,pc(1)).Theoutputfileisallin lowercaseandpackedintolinesof72charactersorless,withtheonlyconcessiontoreadability beingtheterminationoflinesatsemicolonswhenthiscanbedoneconveniently. TheWEBlanguageallowsyoutoprepareasingledocumentcontainingalltheinformationthatis neededbothtoproduceacompilablePascalprogramandtoproduceawell-formatteddocument describingtheprograminasmuchdetailasthewritermaydesire.TheuserofWEBmustbe familiarwithbothTEXandPascal.WEBalsoprovidesarelativelysimple,althoughadequate, macrofacilitythatpermitsaPascalprogramtobewritteninsmalleasily-understoodmodules. Thecommandlineshouldhaveeitheroneortwonamesonit.ThefirstistakenastheWEBfile (and.webisaddedifthereisnoextension).Ifthereisanothername,itisachangefile(and.ch isaddedifthereisnoextension).ThechangefileoverridespartsoftheWEBfile,asdescribedin theWEBsystemdocumentation. TheoutputfilesareaPascalfileandastringpoolfile,whosenamesareformedbyadding.pand .poolrespectivelytotherootoftheWEBfilename. SEEALSO pc(1),pxp(1)(forformattingtangleoutputwhendebugging),tex(1). DonaldE.Knuth,TheWEBSystemofStructuredDocumentation. DonaldE.Knuth,LiterateProgramming,ComputerJournal27,97−111,1984. WayneSewell,WeavingaProgram,VanNostrandReinhold,1989,ISBN0-442-31946-0. DonaldE.Knuth,TEX:TheProgram(VolumeBofComputersandTypesetting),Addison-Wes-ley,1986,ISBN0-201-13437-3. DonaldE.Knuth,METAFONT:TheProgram(VolumeDofComputersandTypesetting),Addison-Wesley,1986,ISBN0-201-13438-1. TheselasttwoarebyfarthelargestextantexamplesofWEBprograms. ThereisanactiveInternetelectronicmaildiscussionlistonthesubjectofliterateprogramming; sendasubscriptionrequesttolitprog-request@shsu.edutojoin. AUTHORS WEBwasdesignedbyDonaldE.Knuth,basedonanearliersystemcalledDOC(implementedby IgnacioZabala).ThetangleandweaveprogramsarethemselveswritteninWEB.Thesystem wasoriginallyportedtoUnixatStanfordbyHowardTrickey,andatCornellbyPavelCurtis. Web2C7.32x 12February1993 1